A sample of 𝑛=20

n

=

20

 observations is drawn from a normal population with 𝜇=950

μ

=

950

 and 𝜎=250

σ

=

250

. Find each of the following:

A. 𝑃(𝑋

¯

>1061)

P

(

X

¯

>

1061

)

Probability = 

B. 𝑃(𝑋

¯

<838)

P

(

X

¯

<

838

)

Probability = 

C. 𝑃(𝑋

¯

>894)

P

(

X

¯

>

894

)

Probability = 


1
Expert's answer
2021-10-11T10:00:29-0400

A.


"P(X>1061)=1-P(X\leq1061)"

"=1-P(Z\leq\dfrac{1061-950}{250/\sqrt{20}})\approx1-P(Z\leq1.9856)"

"\approx0.023537"


B.


"P(X<838)=P(Z<\dfrac{838-950}{250/\sqrt{20}})"

"\approx P(Z<-2.0035)\approx0.022561"

C.


"P(X>894)=1-P(X\leq894)"

"=1-P(Z\leq\dfrac{894-950}{250/\sqrt{20}})\approx1-P(Z\leq-1.0018)"

"\approx0.841770"